OpenStudy (anonymous):

Solve for both x and y by using the following equations. y^2 + (y + 2) ^2 = 100 x = y + 2

OpenStudy (anonymous):

first eqn is y2+2y-48=0 solving for factors, (y-6)(y+8)=0 y=6,-8 substitute y in x=y+2